OpenStudy (anonymous):

system has an infinite number of solutions, use set-builder notation to write the solution set. If a system has no solution, state this m−2n=16,4m+n=1

OpenStudy (anonymous):

m=16+2n from equation 1 now plug this equation into equation 2 4(16+2n)+n=1 64+8n+n=1 9n=1-64 9n=-63 n=-7 =>m=16+2n=16+2(-7)=16-14=2